Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
3x + 7y = 77
-2x + 7y = 42

Answer by edjones(8007)

3x + 7y = 77
-2x + 7y = 42
--------------subtract
5x=35
.
x=7
y=8
.
Ed